Suppose a ϵ Z. Prove by contradiction that if a2 -2a + 7 is even, then a is even.
1
Expert's answer
2020-10-23T14:07:22-0400

Your question is wrong because if "a^2-2a+7" is even then "a" need not to be even.

Suppose "a=1\in \Z" . Then "a^2-2a+7=1^2-2*1+7=6" is an even number. But "a=1" is not even.